A body of mass 600g moves round a track of area 154cm2 in 4secs. Calculate the total time taken to make expend 0.25KJ of energy


1
Expert's answer
2021-04-01T18:57:50-0400

m = 600 gm = 0.6 Kg


t = 4 sec.


Area of circular track = 154 sq. cm

pi r2Β = 154


"r^2 = \frac{154}{pi}"


"r^2 = \frac{154}{3.14}"


r = 7 cm. = 0.07 m


Distance covered = perimeter of circular track = 2 pi r = 2 x 3.14 x 0.07 = 0.44 m


"Speed = \frac{Dist}{time} 
= \frac{0.44}{4} = 0.11 m/sec."


"Kinetic Energy = \frac{1}{2} m v^2"


KE per unit round in 4 sec. = "\frac{1}{2} m v^2 
= \frac{1}{2} 0.6 x 0.11^2 
= 0.00363 J"


"No. of rounds for 0.25 KJ = \frac{0.25 * 1000}{0.00363}
= 68870.52"


"Time for 68870.52 rounds = \frac{68870.52}{4} 
= 17217.6 seconds."


Final Answer:

Total time taken to make expend 0.25KJ of energy = 17217.6 seconds